OpenStudy (anonymous):

A gas sample contains 16.0 g of CH4, 16.0 g of O2, 16.0 g of SO2, and 33.0 g of CO2. What is the total number of moles of gas in the sample? : A) 2.50 moles B) 3.00 moles C) 2.75 moles D) 2.25 moles E) 4.00 moles

OpenStudy (anonymous):

explanation of answer please

OpenStudy (chmvijay):

hey just covert the each gas molecule into moles using the equation moles of gas =gram of gas /Molecular weight of gas for example let us take CH4 moles of CH4= 16/16 =1mole moles of O2 =16/32=0.5 mole just like above find for other two SO2 and CO2 and add all moles of gases u have worked out total number of moles of gas = moles of CH4+Moles of O2+ moles of SO2 + moles of CO2 = 1+0.5+ +